OpenStudy (anonymous):

Hey, Amistre64, can you help me with a derivative, please?

OpenStudy (amistre64):

maybe..whatcha got?

OpenStudy (anonymous):

\[f(x)=e ^{x }-11e ^{-x}-12x\]

OpenStudy (anonymous):

I need to find the local extreme points and the inflection points, which means I need the first and second derivatives. I do not know what to do with that e. It's not a regular e it's that math e. What's it called in math language?

OpenStudy (anonymous):

Are you there ?

OpenStudy (amistre64):

imhere

OpenStudy (amistre64):

the derivative of e^x = e^x teh derivative of -11e^-x = +e^-x derivative of -12x = 12

OpenStudy (amistre64):

e = oilers number = 2.71828182845905...

OpenStudy (anonymous):

But doesn't that e have a funny latin name or something?

OpenStudy (anonymous):

How about the second derivative?

OpenStudy (amistre64):

its named ofter euler, pronounced oiler

OpenStudy (anonymous):

oh, duh! ; )

OpenStudy (amistre64):

the second derivative is similar except for -12 goes to 0 e^x goes to e^x again and +11e^-x for to -11e^-x

OpenStudy (amistre64):

\[f(x) = e^x -11e^{-x}-12x\] \[f'(x) = e^x + 11e^{-x} -12\] \[f''(x) = e^x -11e^{-x}\]

OpenStudy (anonymous):

So do I find the points of inflection and the extrema the way I would as if they were regular numbers? I just don't know what to do with something that isn't a regular number, like this and delta and pi. They really bother me. I don't mind the x's in algebra, but these are way out there.

OpenStudy (amistre64):

e is a regular number; it just has a special name :)

OpenStudy (amistre64):

but yes, find them the regular way

OpenStudy (anonymous):

ok, so if I am finding the points of interest using e, do I use the number value for it or do I just leave it as e?

OpenStudy (amistre64):

you can approximate it with 2.7183

OpenStudy (anonymous):

would I do that normally with a problem that has e in it? Use the number value, I mean. Is that the way it's supposed to be done? Does that link show stuff about e?

OpenStudy (amistre64):

e is an exact value, so any number form of it is going to be an approximation; then way we solve for this is prolly thru logarithms

OpenStudy (amistre64):

ln(x) -11/ln(x) - 12x maybe?

OpenStudy (amistre64):

ln(x) = e and 1/ln(x) = e^(-x)

OpenStudy (amistre64):

well, ln(y) is a better term than x in that case but thats semantics

OpenStudy (anonymous):

This is making my stomach upset, all this crazy math stuff with these weird alien symbols. It's really scary how much I don't know and to think I am getting an A in calculus! Don't ask me how!

OpenStudy (amistre64):

f′(x)= e^x +11e^(−x) −12 = 0 when e^x +11e^(−x) = 12 i dont think theres an easy way to desribe it

OpenStudy (anonymous):

I took a calculus test today and this was on it, finding the extrema and points of inflection, and I just never saw such a problem. It wasn't in our homework problems, and since i am taking this online it's not like I can go ask the instructor. I just don't know what to do with it.

OpenStudy (amistre64):

e^x^2 -12e^x +11= 0 is a quadratic

OpenStudy (amistre64):

do you know what inflection and min and max are?

OpenStudy (amistre64):

a^2 -12a + 11 = 0 when.... (a-11)(a-1) = a = 1 and a = 11

OpenStudy (anonymous):

We just got done with the chapters and I did really well with the trig functions and the regualr numbers in finding the critical points and local max and min and extreme max and min, but this one i didn't work before and i don't have the common sense to be able to apply the regular rules to something like this. If someone doesn't show me step by step, I will not understand. Got a 12.7 out of 15 points on the test.

OpenStudy (amistre64):

when e^x = 1 or 11 we should test this

OpenStudy (amistre64):

that doesnt seem to work out... more than likely it just a matter of using what you know to figure out what you dont think you know

OpenStudy (amistre64):

and an 84% is a B so :)

OpenStudy (anonymous):

But what if you don't know anything? What then? That's the way I feel about calculus most of the time! I work on Math XL sometimes 14 hours a day, and only got a B. My overall greade is a 96% but I can feel it slowly slipping away... How do I find out how to solve that e problem for the critical points and stuff? Any suggestions?

OpenStudy (amistre64):

first derive it; and see what you can do to it algebraicly f'(x) = e^x +11e^(−x) −12 ; we know that a negative exponent is a reciprocal right?

OpenStudy (anonymous):

uh...I guess. Exponents aren't my strong point. None of it really, it's been way too long to just remember!

OpenStudy (amistre64):

normally a negative exponent means to invert it; let me recheck that thought tho

OpenStudy (anonymous):

I doknow that x^-3, say, is 1/x^3

OpenStudy (amistre64):

spose we saw a normal number like 2^x; how would it graph look?

OpenStudy (anonymous):

I don't know.

OpenStudy (amistre64):

if we plot it it looks like this...

OpenStudy (amistre64):

if we plot 2^(-x) it loks like this, the one in red

OpenStudy (anonymous):

dont' mean to be rude, but there wasn't a line in red in that second attachment.

OpenStudy (amistre64):

which is the same graph as (1/2)^x lol.... forgot to save before i attached

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%281%2F2%29^x

OpenStudy (anonymous):

no red line there either. I'm sorry to be such a pain. Just when I think I may actually be not so dumb, I realize I was mistaken. I know nothing about any of this and i really don't know what to do!

OpenStudy (amistre64):

soo... 2^-x is the same as (1/2)^x

OpenStudy (amistre64):

the last jpg has both lines

OpenStudy (amistre64):

e^-x = (1/e)^x ; its just a number like 2 or 5 just written in a way we can express it exactly

OpenStudy (anonymous):

got that graph. It is symmetrical about the y axis. That's about all I know.

OpenStudy (anonymous):

So i might be able to do that e problem with using the value for it? Do I still need logs for that?

OpenStudy (amistre64):

\[e^x +\frac{11}{e^x}-12=0\]

OpenStudy (amistre64):

lets see :)

OpenStudy (amistre64):

\[e^x +\frac{11}{e^x} = 12\] right?

OpenStudy (anonymous):

yes, tht is algebra! i like algebra! alot!

OpenStudy (amistre64):

\[\frac{(e^x)^2 + 11}{e^x} = 12\]

OpenStudy (amistre64):

\[(e^x)^2 + 11 = 12e^x\]

OpenStudy (amistre64):

\[(e^x)^2 -12(e^x) + 11 = 0\]

OpenStudy (amistre64):

u = e^x \[u^2 -12u +11 = 0\]

OpenStudy (amistre64):

solve for u

OpenStudy (anonymous):

don't mean to sound obtuse, but this is the derivative you are solving for 0, right?

OpenStudy (amistre64):

yes :)

OpenStudy (anonymous):

ok that factors to u-11 and u-1, which give u=11 and u=1. Are those critical points, then?

OpenStudy (amistre64):

are we looking to solve an equation with u? or e^x?

OpenStudy (anonymous):

The e one. I need to find critical points so I can find the inflection points and the extrema.

OpenStudy (amistre64):

u = 11, or 1 u = e^x e^x = 11, or 1

OpenStudy (amistre64):

e^x = 11 ; ln(11) = x e^x = 1 ; ln(1) = x now this is calculator time....

OpenStudy (amistre64):

ln(11) = 2.398.... ln(1) = 0 are our critical points

OpenStudy (anonymous):

Does this problem have a domain of negative to positive infinity?

OpenStudy (amistre64):

teh tree terms all have domains that are form -inf to inf

OpenStudy (amistre64):

so yes

OpenStudy (amistre64):

the inflection point is the axis of symettry of the quadratic... so at x = halfway between 0 and 2.398.... is your inflection point

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=%28e^x%29-%2811%2F%28e^x%29%29-12x tells me im right :)

OpenStudy (amistre64):

the main point here is that you see how I turned this into a quadratic, something that you shuold be familiar with alrady

OpenStudy (anonymous):

But do I have to have exact numbers? I mean the way we have been doing this, every point of inflection and max and min all have very specific cooridnates. I'm sorry I' m so dumb at this. I'll bet you're a male, right? Pardon me, but I cannot tell for sure by your user name and I personally think men are much better with numbers. All my math teachers were male, all though my years of school.

OpenStudy (amistre64):

my kids call me 'daddy' :)

OpenStudy (anonymous):

aw, how many?

OpenStudy (amistre64):

2... 13 and 15

OpenStudy (anonymous):

I have a 13 and 15 year old too! I atually have 5, 6,8,11,13,15. 4 girls and a boy. Your kids are so lucky to have such a smart dad!

OpenStudy (amistre64):

:) thnx. Lotta kids on your end, nothing better to do in the winter? ;)

OpenStudy (anonymous):

lol~!!!!!!!!!!!! funny guy! Never wanted kids, actually, but now...they are somethin aren't they? Actually, they will be home in a couple of minutes. Almost summer now, I bet it's HOT there already!

OpenStudy (amistre64):

its been chillly these past few days, had a front move thru and cooled us down

OpenStudy (anonymous):

Cooled you down to what? 83 instead of 85? lol

OpenStudy (amistre64):

lotta kids is is a sign of catholic or mormon... statistically speaking that is

OpenStudy (amistre64):

:) water freezes at 60 here

OpenStudy (anonymous):

my husband is a reformed catholic. We are both lutheran now. But they don't believe in birth control either. Ha!

OpenStudy (amistre64):

I grew up with German Catholic family; joined the Mormons with mom back in 84 tho and aint looked back since :)

OpenStudy (anonymous):

not familiar with the mormon religion. Just took a christian worldview course before this calculus. It was awesome! I know quite a few people in Florida. do you like florida?

OpenStudy (amistre64):

florida is ok..... been here for about 40 years now; borned in Lakeland and lived in dadecity area most me life.... thats north of tampa

OpenStudy (amistre64):

lived in north carolina for a few months after mom died... that was nice country

OpenStudy (amistre64):

the exact form would be ln(11) and 0 for our critical points

OpenStudy (anonymous):

ok and badk to math...sorry. I am definitely female; nosy and all! Ssorry! But how do you know all this math stuff? Were you a teacher ever? Or do you just like numbers?

OpenStudy (amistre64):

\[f''(x)=e^x -\frac{11}{e^x}\]

OpenStudy (amistre64):

ive always liked math; but grew up doing labor and was a carpenter for 15 years, hurt my back and decided to get a degree in Math to teach at college